OpenStudy (anonymous):

could 5x+4=3x^2-2 be solved using the quadratic formula after being rearranged and simplified?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yah.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or -3x^2+9x+3x^2=14+1-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that ain't right

OpenStudy (anonymous):

3x^2-5x-6=0

OpenStudy (anonymous):

-3x^2 + 5x +6